# Connecting Neural Models Latent Geometries with Relative Geodesic Representations

### v1

@@ 3.2 Relative geodesics representations
Figure 2: Pairwise latent-space distance matrices for (a) MNIST and (b) CIFAR-10. In each subfigure, the left heatmap shows the straight-line energy proxy and the right shows the full Riemannian geodesic distances. The Spearman rank correlation between the two measures is $\rho=0.99$ for MNIST and $\rho=1.00$ for CIFAR-10, demonstrating near-perfect agreement.
@@ 3.3 Choice of pullback metric
For autoencoders, it has been argued that a pullback metric is beneficial to reflect the underlying geometry of the latent space (Tosi et al., 2014; Arvanitidis et al., 2018; Hauberg, 2019). For discriminative models, such as classifiers, it is not immediate how to assign a Riemannian structure to the space of latent representations. From the perspective of information geometry, perhaps the most natural choice is the Fisher information matrix (Amari, 2016), in which case the metric in the output space can be obtained as the one with Categorical likelihood. However, neural networks typically experience Neural Collapse (Kothapalli, 2023), rendering the resulting geometry troublesome. Here we discuss two principled approaches: pullback and Diet.
Pullback.   Perhaps the most natural idea is, as discussed in Section 3.1, to construct a pullback metric based on the model’s outputs. In practice, given a model, we train a classification head upon the latent representations and utilize the resulting Riemannian structure.
@@
Diet.   Diet (Ibrahim et al., 2024) is a simple self-supervised training method, based on instance discrimination task, which has been shown to yield identifiability guarantees (Reizinger et al., 2025). Specifically, it can identify the cluster centers of Von-Mises Fisher (VMF) distributions, which lie on a unit sphere.
One can consider such a scenario (Reizinger et al., 2025): some latent variables $z$ are drawn from a VMF distribution, which naturally lie on a unit sphere, and pushed forward through an injective generator function $g$ to obtain the data $x$. Given only $x$ without the knowledge of $g$, it is possible to recover the latent variables $z$ through parameterizing a model and optimizing
$L=\mathbb{E}_{x,i}\left[-\log\frac{\exp\left(\langle w_{i},f(x)\rangle\right)}{\sum_{j}\exp\left(\langle w_{j},f(x)\rangle\right)}\right],$
where $w$ is a linear layer without bias and $f$ is a nonlinear encoder. After the model is trained using the above criterion, up to some assumptions, when both $f$ and $w$ are not unit-normalized, $f\circ g$ is linear. This hints that $f\circ g$ may tell us something about the underlying spherical structure of $z$.
While Diet was proposed to train the entire neural network (Ibrahim et al., 2024), we use it to learn a classification head on top of the pretrained neural network. Specifically, we add several layers on top of the pretrained model, and use it as $f$. After the model is trained, we may expect $f\circ g$ to give the ground truth $z$ up to a linear transformation. Furthermore, Lähner and Moeller (2024a) noted that it was beneficial to employ data augmentations, which we also observed to be important to achieve good performance.
We propose to directly employ the geodesic distances of a sphere to form the relative representations in Diet. Specifically, on a unit sphere, the distance between two points $x$ and $y$ is given by
$d(x,y)=\arccos\left(\frac{x^{\top}y}{\left\lVert x\right\rVert\left\lVert y\right\rVert}\right),$
which interestingly bears a strong resemblance to the cosine distances as used in the original paper on relative representations (Moschella et al., 2023). Observe that $x$ and $y$ do not need to lie precisely on the unit sphere; instead, they are projected onto the unit sphere. As such, this implies the assumption that different instances of $f\circ g$ yield transformations that are constant scalings, which is a stronger assumption than a linear transformation. Moreover, that the data lies on a unit sphere is a rather strong assumption. Nevertheless, as will be shown later, this approach results in meaningful representations across the models.
4 Experimental evaluation on Autoencoders
@@
In the following, we evaluate relative geodesic representations on the latent communication problem across models trained with different initializations, different architectures, and tasks.
4.1 Aligning neural representational spaces trained independently
Figure 3: Aligning latent spaces of autoencoders: MRR score as a function of the number of anchors on pairs of autoencoders trained with different initializations on the MNIST (left), FashionMNIST (center), CIFAR10 (right) datasets, respectively. In green, we plot the performance of Moschella et al. (2023), in blue, our method. The shaded area indicates standard deviation across different random sets of anchors. Relative geodesics consistently outperform the cosine baseline, obtaining peak performance.
Experimental setting.   For the following experiment we trained pairs of convolutional autoencoders $(F_{1},F_{2})$ with different initializations on the MNIST Deng (2012), FashionMNIST Xiao et al. (2017), CIFAR10 Krizhevsky (2009) datasets. The architecture of the convolutional autoencoder is detailed in the Appendix. After training, we extracted 10k samples from the test set, and mapped them to the latent spaces of the two models, to representations $\mathbf{Z}_{1}=E_{1}(\mathbf{X}),\mathbf{Z}_{2}=E_{2}(\mathbf{X})$ respectively. Starting from a small set of anchors in correspondence $\mathcal{A}_{\mathcal{X}}\mapsto\mathcal{A}_{\mathcal{Y}}$, the objective is to evaluate how well it is possible to recover the full correspondence between the representations $\mathbf{Z}_{1},\mathbf{Z}_{2}$ from the relative representations. As a baseline, we compare with relative representations using cosine similarity Moschella et al. (2023).
Analysis of results.   Fig. 3 plots the performance in terms of MRR on MNIST, FashionMNIST, CIFAR10 datasets. To obtain the score we first compute similarity matrices between relative representations of the two spaces as $\mathbf{D}(\mathbf{Z}_{1},\mathbf{Z}_{2})$ where $\mathbf{D}_{i,j}=\frac{RR(\mathbf{Z}_{1})_{i}^{T}RR(\mathbf{Z}_{2})_{j}}{\|RR(\mathbf{Z}_{1})_{i}\|_{2}\|RR(\mathbf{Z}_{2})_{j}\|_{2}}$. Then we compute the Mean Reciprocal Rank (MRR, see Appendix A.1.1) on top of the similarity matrix. In the figure, we plot MRR as a function of a random set of anchors, where the shaded areas indicate the standard deviations over 5 different sets of random anchors with the same cardinality. Our method consistently performs better than Relative Representation, saturating the score with few anchors on all the domains, despite the different degrees of complexity of the latent spaces. In addition, our method shows significantly less variance, being more robust to the choice of the anchor set.
Takeaway: Relative geodesic representation near-perfectly captures transformations between representational spaces of models initialized differently, outperforming Moschella et al. (2023) in sample efficiency and robustness.
4.2 Stitching autoencoder models
Figure 4: Stitching on Autoencoders: We visualize qualitative reconstructions of samples, stitching autoencoders of models trained with different initializations on MNIST (left), FashionMNIST (center), CIFAR10 (right). The first two columns show reconstructions from the original models; the middle columns represent baselines Maiorca et al. (2024); Lähner and Moeller (2024a); Moschella et al. (2023); the rightmost column is our method. Relative geodesics yield the best stitching results using just 5 anchors.
Experimental setting.   We consider the same pairs of autoencoders trained on the MNIST, FashionMNIST, CIFAR10 datasets of section 3. Starting from a set of five random anchors, we estimate a transformation $T$ between the model representational spaces $Z_{1},Z_{2}$. Different from Moschella et al. (2023), in which zero-shot stitching was achieved by training once a decoder module with relative representations and then exchanging different encoder modules, here we achieve stitching without training any decoder. We compute relative representation with respect to the set of anchors, and compute a similarity matrix $\mathbf{D}(\mathbf{Z}_{1},\mathbf{Z}_{2})$. Then we compute the vector $\mathbf{c}=\operatorname*{arg\,max}_{i}(\mathbf{D})$ representing a correspondence between the two representations matrices $\mathbf{Z}_{1}$, $\mathbf{Z}_{2}$, and use $c$ to fit a linear transformation $T$ to approximate the transformation between the two domains. We perform stitching by performing the following operation for a sample $x\in\mathcal{X}$: $\tilde{x}=D_{2}\circ T\circ E_{1}(x)$.
Analysis of results.   We visualize the results of reconstructions of random samples in Fig. 4, comparing with Moschella et al. (2023); Lähner and Moeller (2024a); Maiorca et al. (2024). For each dataset, each column represents respectively: (i) the original autoencoding mapping for a sample $x$ of model $F_{1}$, $D_{1}(E_{1}(x))$, (ii) $D_{2}(E_{2}(x))$, (iii) the mapping $D_{2}(E_{1}(x))$, (iv) the mapping $D_{2}(T_{anchors}E_{1}(x))$ where $T_{anchors}$ is estimated on the five available anchors, (v) the mapping $D_{2}(T_{cosine}E_{1}(x))$ where $T_{cosine}$ is estimated among all 10k samples with the correspondence $c$ obtaining in the relative space of Moschella et al. (2023), (vi) Our result $D_{2}(T_{relgeo}E_{1}(x))$, where $T_{relgeo}$ is estimated from the correspondence obtained in the relative geodesic space. While the baselines do not reach a good enough reconstruction quality, reconstructions with our method are almost perfect in accordance with the results in Fig. 3.
Takeaway: The relative geodesic space enables stitching neural modules trained on different seeds.
5 Experiments on vision foundation models
@@
In this section we evaluate relative geodesic representations performance on retrieval and model stitching tasks on vision foundation discriminative models across models trained with different objectives, architectures, and sizes.
5.1 Matching representational spaces of discriminative foundation models
@@
Rel(Cosine) Moschella et al. (2023) $0.129\pm 0.135$ $0.166\pm 0.162$ $0.221\pm 0.178$ $0.135\pm 0.148$ $0.068\pm 0.08$
RelGeo(Pullback) $0.047\pm 0.013$ $0.112\pm 0.031$ $0.412\pm 0.09$ $0.28\pm 0.129$ $0.025\pm 0.012$
@@
In this section, we test the compatibilities of representations of vision foundation models with different architectures, such as residual networks He et al. (2016b) and vision transformers Dosovitskiy et al. (2021), and with different pretraining objectives including classification and self-supervised learning.
Experimental setting.   We perform experiments on retrieval tasks on pretrained vision foundation models, investigating how well we can match representations together with different backbones subject to the decoding tasks, on 5 datasets, varying in complexity and size: CIFAR10, CIFAR100 Krizhevsky (2009), SVHN Yuval Netzer et al. (2011), CUB Wah et al. (2023), and ImageNet-1k Russakovsky et al. (2015). For ImageNet-1k, we used $1000$ anchors, while for other datasets we used $500$. As backbones we consider ResNet-50 (He et al., 2016a), Vision Transformers (ViT) (Dosovitskiy et al., 2021), with both patch 16-224 and patch 32-384, and DINOV2 (Oquab et al., 2024). We compare the original formulation of relative representations with cosine similarity Moschella et al. (2023) denoted as Rel(Cosine), relative geodesic representation using Euclidean pullback metric denoted as RelGeo(Pullback), and pulling back the spherical metric using a Diet decoder denoted RelGeo(Diet).
Analysis of results.   Table 1 shows results from different methods averaged across all possible pairs of models on the considered datasets. Additionally, Fig. 5 shows the results on CUB datasets, highlighting the numbers in different settings. While RelGeo(Pullback) may result in worse MRR numbers, RelGeo(Diet) provides consistently improved retrieval performance. In the Appendix we report full results for every dataset.
Takeaway: Relative geodesic representations pulling back from instance discrimination decoders are identifiable across vision foundation models, improving retrieval performances.
Figure 5: CUB Accuracies (top) and symmetricized MRR cosine (bottom)
5.2 Zero-shot stitching of vision foundation models
Table 2: Average stitching performances across different settings. RelGeo(Pullback) often outperforms Rel(Cosine), while RelGeo(Diet) remains competitive.

### v2

@@ 3.2 Relative geodesics representations
Figure 2: Pairwise latent‐space energy matrices for (a) MNIST and (b) CIFAR-10. In each subfigure, the left heatmap shows the straight-line energy approximation and the right shows the geodesic energies of the ground truth geodesic curve. The Spearman rank correlations between the two measures are $\rho=0.99$ for MNIST and $\rho=1.00$ for CIFAR-10, demonstrating near-perfect agreements.
@@ 3.3 Choice of pullback metric
The properties of the relative geodesic representations are determined by (i) the choice of the output space, (ii) the choice of the metric to pullback from the output space and (iii) the pretraining objective (e.g. reconstruction or classification) on which the decoder was trained.
Generative models.  For models trained on a reconstruction loss such as autoencoders, or on generative objectives, such as variational autoencoders Kingma and Welling (2013), pulling back metrics such as $L2$ distance have been shown to effectively reflect the underlying geometry of the latent space (Tosi et al., 2014; Arvanitidis et al., 2018; Hauberg, 2019).
@@
Discriminative models. For discriminative models, such as classifiers or instance based discriminative models (Ibrahim et al., 2024), it is not immediate how to assign a Riemannian structure to the space of latent representations. From the perspective of information geometry, perhaps the most natural choice is the Fisher information matrix (Amari, 2016), in which case the metric in the output space can be obtained as the one with categorical likelihood. However, neural networks typically experience Neural Collapse (Kothapalli, 2023), possibly rendering the resulting geometry troublesome. We empirically inspect this approach in Appendix A.4.11. In this work we consider two principled approaches for discriminative models based on classification decoder heads and instance discrimination heads.
Pulling back from classifiers.  Perhaps the most natural idea is, as discussed in Section 3.1, to construct a pullback metric based on the model’s outputs, by simply pulling back the euclidean $L2$ metric from the logit space of the classifier. Given an arbitrary encoder model, we train a classification head upon the latent representations (extracted e.g. from the last layer) and pulling back the euclidean $L2$ metric from the output logits of the head. The resulting relative geodesics representation will inherit properties of both the decoder head (up to class information) and the pretrained encoder.
@@
Pulling back from instance discrimination decoders.  Diet (Ibrahim et al., 2024) is a self-supervised training method which has been shown to learn representations with strong generalization to downstream tasks, and yield identifiability guarantees (Reizinger et al., 2025). Specifically, in the infinite data limit representations from the Diet objective align the cluster centers of von-Mises Fisher (vMF) distributions, which lie on a unit sphere. The loss based on simple instance discrimination is:
$\mathcal{L}_{Diet}=\mathbb{E}_{\bm{x},i}\left[-\log\frac{\exp\left(\bm{w}_{i}^{\top}f(\bm{x})\right)}{\sum_{j}\exp\left(\bm{w}_{j}^{\top}f(\bm{x})\right)}\right],$ (6)
where $\bm{W}$ is a linear projection and $f$ is a nonlinear map. Furthermore, assigning the same instance label to data augmentations was shown beneficial to improve invariance. While it was originally proposed to train the entire neural network (Ibrahim et al., 2024), we instead use it to learn a decoder $D$ on top of the pretrained neural network latent representations, by setting $D=f\circ\bm{W}$.To construct relative geodesic representations, we propose to pullback the spherical metric from the penultimate layer of the diet decoder, before the projection head $\bm{W}$. Further discussions on Diet can be found in Appendix A.2.1.
Both classifier and instance discriminator approaches discussed above use proper pullback metrics and fall under our proposed framework of relative geodesic representations: these representations inherit semantic information, up to class or instance level, from the decoder, while retaining structure of the pretrained encoder. Notably, when the encoder is pretrained, the relative geodesics representations are computationally efficient, since the decoder remains lightweight and the approximate geodesic energy computation is cheap. As we demonstrate in the following experimental sections, this approach yields meaningful and identifiable representations that are consistent across models.
4 Experiments
@@
In the following sections we will evaluate the performance of relative geodesic representations on two instances of the latent communication problem, across models with different initializations, architectures, sizes and modalities.
Tasks description.  We evaluate our approach on two representative instantiations of the latent communication problem: retrieval and neural stitching. In a retrieval setting we aim to solve the latent communication problem up to the instance level. Given pairs of model encoders $(E_{\mathcal{X}}^{1},E_{\mathcal{X}^{\prime}}^{2})$ and access to their latent representations, we seek to recover a full correspondence $\Lambda$ starting from a partial one $\Gamma$. For neural stitching the goal is to solve the latent communication problem up to the task-label level. Classical stitching approaches train an adapter $\Psi$ between intermediate components of distinct neural networks so that $D_{\mathcal{X}^{\prime}}^{2}\circ\Psi\circ E_{\mathcal{X}}^{1}$ remains functional on a downstream task (e.g., classification). In Section 4.2, we operate in the zero-shot stitching regime (Moschella et al., 2023), where no adapter is trained explicitly. Instead, we solve implicitly for the transformations between representations by mapping them into relative representation spaces. This enables stitching pairs of models without any fine-tuning or additional supervision.
@@
In Section 4.1, we evaluate relative geodesic representations on generative models, focusing on autoencoders. This analysis examines performance across networks trained with different initializations and datasets. In Section 4.2, we extend the evaluation to discriminative foundation models, assessing performance at scale across diverse architectures, pretraining objectives (e.g., self-supervised and classification), datasets, and modalities.
4.1 Experimental evaluation on autoencoders
In the following sections, we evaluate relative geodesic representations on the latent communication problem across autoencoder models trained with different initializations, architectures and datasets.
4.1.1 Aligning independently trained neural representational spaces
Figure 3: Aligning latent spaces of autoencoders: MRR score as a function of the number of anchors on pairs of autoencoders trained with different initializations on the MNIST (left), FashionMNIST (center), CIFAR10 (right) datasets, respectively. In green, we plot the performance of Moschella et al. (2023); in red and orange the linear and orthogonal baselines respectively; in blue, our method. The shaded area indicates standard deviation across 5 different random sets of anchors. Relative geodesic consistently outperforms baselines, obtaining peak performance.
Setting.  For the following experiment, we trained pairs of convolutional autoencoders $(F_{1},F_{2})$ with different initializations on MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), CIFAR10 (Krizhevsky, 2009) datasets. The architecture of the convolutional autoencoder is detailed in Appendix A.3.2. After training, we extracted 10k samples from the test set, and mapped them to the latent spaces of the two models, to representations $\mathbf{Z}_{1}=E_{1}(\mathbf{X}),\mathbf{Z}_{2}=E_{2}(\mathbf{X})$ respectively. Starting from a small set of anchors in correspondence $\Gamma:\mathcal{A}_{\mathcal{X}}\mapsto\mathcal{A}_{\mathcal{Y}}$, the objective is to evaluate how well it is possible to recover the full correspondence $\Lambda$ between the representations $\mathbf{Z}_{1},\mathbf{Z}_{2}$ from the relative representations. As a baseline, we compare with relative representations using cosine similarity (Moschella et al., 2023), and with fitting a linear or orthogonal mapping using $\Gamma$.
Analysis of results.  Fig. 3 plots the performance in terms of MRR on MNIST, FashionMNIST and CIFAR10 datasets. To obtain the score, we first compute similarity matrices between relative representations of the two spaces as $\mathbf{D}(\mathbf{Z}_{1},\mathbf{Z}_{2})$ where $\mathbf{D}_{i,j}=\frac{RR(\mathbf{Z}_{1})_{i}^{T}RR(\mathbf{Z}_{2})_{j}}{\|RR(\mathbf{Z}_{1})_{i}\|_{2}\|RR(\mathbf{Z}_{2})_{j}\|_{2}}$. Then we compute the Mean Reciprocal Rank (MRR, see Appendix A.3.1) on top of the similarity matrix. In the figure, we plot MRR as a function of a random set of anchors, where the shaded areas indicate the standard deviations over 5 different sets of random anchors with the same cardinality. Our method consistently performs better than Moschella et al. (2023), saturating the score with few anchors on all the domains, despite the different degrees of complexity of the latent spaces. In addition, our method shows significantly less variance, being more robust to the choice of the anchor set.
Takeaway. Relative geodesic representation near-perfectly captures transformations between representational spaces of models initialized differently, sample efficiency and robustness.
4.1.2 Stitching autoencoder models
[…]
Figure 4: Stitching on Autoencoders: We visualize qualitative reconstructions of samples, stitching autoencoders of models trained with different initializations on MNIST (left), FashionMNIST (center), CIFAR10 (right). The first two columns show reconstructions from the original models; middle three columns represent baselines (Maiorca et al., 2024; Moschella et al., 2023); the rightmost column is our method. Relative geodesic yields the best stitching results using just 5 anchors.
Setting.  We consider the same pairs of autoencoders trained on the MNIST, FashionMNIST, CIFAR10 datasets of Section 3. Starting from a set of five random anchors, we estimate a transformation $T$ between the model representational spaces $Z_{1},Z_{2}$. In this experiment, to keep differently from Moschella et al. (2023), in which zero-shot stitching was achieved by training once a decoder module with relative representations and then exchanging different encoder modules, here we achieve stitching without training any decoder. We compute relative representations with respect to the set of anchors, and compute a similarity matrix $\mathbf{D}(\mathbf{Z}_{1},\mathbf{Z}_{2})$. Then we compute the vector $\mathbf{c}=\operatorname*{arg\,max}_{i}(\mathbf{D})$ representing a correspondence between the two representation matrices $\mathbf{Z}_{1}$, $\mathbf{Z}_{2}$, and use $c$ to fit a linear transformation $T$ to approximate the transformation between the two domains. We perform stitching by performing the following operation for a sample $x\in\mathcal{X}$: $\tilde{x}=D_{2}\circ T\circ E_{1}(x)$.
Analysis of results.  We visualize the results of reconstructions of random samples in Fig. 4, comparing against Moschella et al. (2023); Lähner and Moeller (2024); Maiorca et al. (2024). For each dataset, each column represents respectively: (i) the original autoencoding mapping for a sample $x$ of model $F_{1}$, $D_{1}(E_{1}(x))$, (ii) $D_{2}(E_{2}(x))$, (iii) the mapping $D_{2}(E_{1}(x))$, (iv) the mapping $D_{2}(T_{anchors}E_{1}(x))$ where $T_{anchors}$ is estimated on the five available anchors, (v) the mapping $D_{2}(T_{cosine}E_{1}(x))$ where $T_{cosine}$ is estimated among all 10k samples with the correspondence $c$ obtaining in the relative space of Moschella et al. (2023), (vi) Our result $D_{2}(T_{relgeo}E_{1}(x))$, where $T_{relgeo}$ is estimated from the correspondence obtained in the relative geodesic space. As shown in Fig. 4, while the baselines do not reach a good enough reconstruction quality, reconstructions with our method are almost perfect in accordance with the results in Fig. 3.
Takeaway. Relative geodesics enable stitching of neural modules trained with different initializations.
4.2 Experiments on vision foundation models
In this section, we evaluate relative geodesic representations’ performances on retrieval and model stitching tasks on vision foundation discriminative models across models pretrained with different objectives, architectures, sizes and modalities.
4.2.1 Matching representational spaces of discriminative foundation models
@@
Rel(Cosine) (Moschella et al., 2023) $0.129\pm 0.135$ $0.166\pm 0.162$ $0.221\pm 0.178$ $135\pm 0.148 0.0$$068\pm 0.08$
RelGeo($L2$) $0.047\pm 0.013$ $0.112\pm 0.031$ $0.412\pm 0.09$ $28\pm 0.129 0.0$$025\pm 0.012$
@@
In this section, we test the compatibilities of representations of vision foundation models with different architectures, such as residual networks (He et al., 2016) and vision transformers (Dosovitskiy et al., 2021), and with different pretraining objectives including classification and self-supervised learning.
Setting.  We perform experiments on retrieval tasks on pretrained vision foundation models, investigating how well we can match representations together with different backbones subject to the decoding tasks, on 5 datasets, varying in complexity and size: CIFAR10, CIFAR100 (Krizhevsky, 2009), SVHN (Yuval Netzer et al., 2011), CUB (Wah et al., 2023), and ImageNet-1k (Russakovsky et al., 2015). For ImageNet-1k, we used $1000$ anchors, while for other datasets we used $500$. As backbones we consider ResNet-50 (He et al., 2016), Vision Transformers (ViT) (Dosovitskiy et al., 2021) with both patch 16-224 and patch 32-384, and DINOv2 (Oquab et al., 2024). We compare the original formulation of relative representations with cosine similarity (Moschella et al., 2023) denoted as Rel(Cosine), relative geodesic representations pulling back from Euclidean logits denoted as RelGeo($L2$), and pulling back the spherical metric using a Diet decoder denoted as RelGeo(Diet).
Analysis of results.  Table 1 shows results from different methods averaged across all possible pairs of models on the considered datasets. Additionally, Fig. 5 shows the results on CUB. While RelGeo($L2$) may result in worse MRR numbers, RelGeo(Diet) provides consistently improved retrieval performance. In Appendix A.4.1 we report full results for the datasets.
Takeaway. Relative geodesic representations pulling back from instance discrimination decoders are identifiable across vision foundation models, improving retrieval performances.
Figure 5: CUB Accuracies (top) and symmetricized MRR cosine (bottom). RelGeo(Diet) and especially RelGeo($L2$) provide strong stitching accuracies, while RelGeo(Diet) maintains strong instance identifiability.
4.2.2 Zero-shot stitching of vision foundation models
Table 2: Average stitching performances across different settings. RelGeo($L2$) often outperforms Rel(Cosine), while RelGeo(Diet) remains competitive.